OpenStudy (anonymous):

i kind of know its C need 2nd opinion A tire manufacturer took a random sample of 50 tires and found that 3 were defective. In a shipment of 2400 tires, how many could be expected to have defects? A. 6 B. 144 C. 150 D. 2353

OpenStudy (danjs):

The same ratio...scaled up \[\frac{ 3 }{ 50 } = \frac{ x }{ 2400 }\]

OpenStudy (anonymous):

so is it C or not?

OpenStudy (anonymous):

@DanJS

OpenStudy (anonymous):

the greatest help in the world

OpenStudy (danjs):

cross multiply 2400 * 3 = 50 *x

OpenStudy (danjs):

divide by 50 x = 144

OpenStudy (danjs):

@speartonion